### v1

@@ 5.2 Travel Demand Prediction
As shown in Fig. 7, BNF produces accurate travel demand forecasts with well-calibrated 95% prediction intervals at selected Chicago pickup locations. The model effectively captures both the temporal dynamics of demand and the underlying predictive uncertainty.
@@ 5.3 Nested Benders Decomposition Performance
Table 5 reports the runtime, number of iterations, and the relative optimality gaps across three scenario set sizes. For the smallest instance (135 scenarios), NBD converges within 10 iterations and achieves a near-exact solution with a final gap of 0.08%. Increasing the number of scenarios to 540 raises the runtime to about 90 seconds, but convergence is still reached in fewer than 10 iterations, with a very small median gap of 0.09%.
For the largest instance (1035 scenarios), the runtime increases to nearly 6 minutes and the number of iterations to 15. While the maximum gap observed during the iterations is higher (4.52%), the algorithm ultimately converges to a high-quality solution with a median gap of 1.50%. These results show that runtime grows with the number of scenarios, but NBD remains tractable and consistently produces solutions with small gaps even for problem sizes where the extensive form would be unsolvable.
Table 5: Runtime and convergence diagnostics across scenario counts.
@@ 5.4 EAMoD Optimization Performance Comparison
Energy costs and V2G revenues show similar trends in both cities. Under flat tariffs, all methods incur costs of around €0.06/kWh with negligible revenue. However, under peak and solar tariffs, SMPC and DMP consistently achieve the lowest effective energy costs and the highest revenues. In Chicago, SMPC earns almost €975 per vehicle under solar pricing, which is approximately double the value in San Francisco of €509, reflecting the larger fleet and heavier charging activity. RCC captures somewhat less revenue due to its conservative scheduling, while RB never participates in V2G and pays the highest electricity prices.
@@
Table 7: Average electricity price paid per kilowatt-hour (€/kWh) and corresponding vehicle-to-grid (V2G) revenue (€) under flat, peak, and solar pricing schemes for Chicago and San Francisco.
@@ 5.6 Sensitivity analysis of battery size and energy efficiency (San Francisco)
Vehicles 2 and 3 display greater sensitivity to battery size. For Vehicle 2, the median wait time decreases slightly from 225s at 40kWh to 217s at 70 kWh, while the share of passengers waiting more than 600s falls from 154 to 151. Expanding to 100 kWh provides only marginal change: median 219 s (slightly higher), with >600 s falling to 144. This suggests that a 70 kWh battery is sufficient to mitigate charging as a performance bottleneck for Vehicle 2, with further increases offering diminishing returns. For Vehicle 3, the least efficient option, the effect of battery size is far more pronounced. With 40 kWh, the system exhibits very poor performance, with a median wait time of 325s, a 95th percentile of 2625s, and more than 3600 passengers experiencing delays over 600s. Upgrading to 70 kWh brings dramatic improvements: the median drops to 219s, the 95th percentile to 432s, and delays over 600s fall by more than 95% to just 145. Increasing to 100 kWh does not significantly improve service further, with a median of 222s and 149 delays over 600s. This highlights that for inefficient vehicles, battery size is a critical driver of feasibility, with performance plateauing around 70 kWh.
@@
These findings are consistently reflected in other system-level metrics. Figures 9 demonstrate that increasing fleet size and charger availability effectively reduce wait times, particularly for the less efficient Vehicles 2 and 3. Vehicle 1, however, continues to exhibit stable performance, characterized by lower 75th percentile and minimal sensitivity to infrastructure constraints. Furthermore, Vehicle 3 has the highest energy consumption and consequently the greatest charging demand, especially during peak hours, see Fig. 10(a). Conversely, Vehicle 1 maintains a smoother, more aligned charging-consumption profile throughout the day.
Fig.11(a) reports the estimated number of daily charge cycles (defined as charging from 20% to 80% state of charge) across vehicle types and battery sizes. As anticipated, larger batteries consistently require fewer charging cycles. Vehicle 1 consistently exhibits the lowest and most stable charging frequency, requiring just over one cycle per day even with a 40 kWh battery. In contrast, Vehicles 2 and 3 demonstrate higher and more variable charging frequencies, particularly with smaller battery configurations. For instance, Vehicle 3 with a 40 kWh pack exhibits a median of 1.5 cycles per day and substantial variability in its charging behavior. Increasing the battery capacity to 70 kWh or 100 kWh nearly halves the median number of cycles and significantly narrows the distribution of charging events.
@@ 6 Conclusion
This paper introduced a combined stochastic and robust optimization framework for Electric Autonomous Mobility-on-Demand (EAMoD) operations under multiple uncertainties. By integrating Bayesian neural networks for demand and charger availability forecasting, clustering-based scenario tree reduction, and a tailored Nested Benders Decomposition algorithm, we developed a scalable stochastic model predictive control (SMPC) approach capable of handling temporally correlated, multivariate uncertainties. High-fidelity simulations in San Francisco and Chicago demonstrate that SMPC reduces median passenger wait times by up to 36%, 95th percentile wait times by nearly 20%, rebalancing distance by up to 27%, and electricity costs by more than 35% relative to established baselines. These gains are consistent across two cities with very different demand intensities, spatial layouts, and fleet sizes, underscoring the framework’s robustness and transferability to diverse urban contexts.
Sensitivity analyses revealed that vehicle efficiency and battery size critically affect system performance. Energy-efficient vehicles maintain stable service levels even with small (40 kWh) batteries, requiring fewer charge cycles and exhibiting minimal sensitivity to infrastructure constraints. In contrast, less efficient vehicles show significant performance degradation with small batteries and benefit most from increases up to 70 kWh, beyond which improvements plateau, especially in high-demand settings such as Chicago. Moreover, efficiency not only drives service quality but also expands opportunities for energy-market participation: efficient vehicles consistently generate higher vehicle-to-grid (V2G) revenues, as their lower energy consumption leaves more capacity for grid interaction. Taken together, the results highlight the importance of jointly optimizing control strategies, vehicle selection, and battery configuration to ensure scalable, reliable, and sustainable EAMoD operations.

### v2

@@ Abstract
{highlights}
Stochastic and robust control model manages demand, energy, time, and charging uncertainties
Bayesian neural networks generate multivariate travel and charging demand scenarios
Scenario reduction via soft clustering preserves key patterns of future uncertainty
Tailored Nested Benders decomposition efficiently solves the large-scale MILP model
Simulation results show reduced costs and improved reliability over baseline methods
@@ 5.2 Travel Demand Prediction
As shown in Fig. 7, BNF generates accurate travel demand forecasts with well-calibrated 95% prediction intervals across selected Chicago pickup locations. While the median prediction does not perfectly match realized demand, the uncertainty estimates provide valuable additional information. This highlights the importance of considering predictive uncertainty in downstream decision-making and motivates the construction of scenario trees based on the probabilistic forecasts.
@@ 5.3 Nested Benders Decomposition Performance
In the nested Benders decomposition, one optimality gap is computed for each master problem and subproblem across the scenario tree. Since these gaps can vary considerably across nodes, we report both the maximum and the median gap as indicators of convergence quality. The maximum gap highlights the worst-case convergence behavior across all nodes, while the median gap reflects the typical convergence performance and is less sensitive to outliers. The optimality gap reported here refers specifically to the Benders decomposition procedures. In each iteration, the Benders cuts provide a lower bound on the subproblem cost for the current master solution, while the incumbent solution provides an upper bound. The relative gap is measured as
$\text{Gap}\;=\;\frac{UB-LB}{UB}\times 100\%,$
where $UB$ denotes the incumbent upper bound and $LB$ the lower bound from the cuts. In the ideal case, this gap equals $0\%$, which means that the decomposition has reached the exact optimum. However, driving the gap fully to zero can be computationally prohibitive. In practice, we terminate the NBD algorithm if the gap does not improve for five consecutive iterations or if the runtime exceeds 600 seconds. The reported gaps reflect only the internal convergence of the decomposition. Since the subproblems are solved in relaxed form, the true duality gap with respect to the extensive formulation remains unobservable without solving the full problem.
Table 5 reports the runtime, number of iterations, and the relative optimality gaps across three scenario set sizes. For the smallest instance (135 scenarios), NBD converges within 10 iterations and achieves a near-exact solution with a final max gap of 0.08%. Increasing the number of scenarios to 540 raises the runtime to about 90 seconds, but convergence is still reached in fewer than 10 iterations, with a small median gap of 0.09%.
For the largest instance (1035 scenarios), the runtime increases to nearly 6 minutes and the number of iterations to 15. While the maximum gap observed during the iterations is higher (4.52%), the algorithm ultimately converges to a high-quality solution with a median gap of 1.50%. These results demonstrate that, although runtime grows with the number of scenarios, NBD remains computationally tractable and consistently delivers solutions with small optimality gaps, even for problem sizes where solving the extensive form is impractical.
Table 5: Runtime, number of iterations, and relative optimality gaps of NBD across different scenario counts.
@@ 5.4 EAMoD Optimization Performance Comparison
Energy costs and V2G revenues show similar trends in both cities. Under flat tariffs, all methods incur €0.06/kWh in costs and generate no revenue, as expected. However, under peak and solar tariffs, SMPC and DMP consistently achieve the lowest effective energy costs and the highest revenues. In Chicago, SMPC earns almost 975 € under solar pricing, which is approximately double the value in San Francisco of €509, reflecting the larger fleet and heavier charging activity. RCC captures somewhat less revenue due to its conservative scheduling, while RB never participates in V2G and pays the highest electricity prices.
@@
Table 7: Average electricity price paid per kilowatt-hour (€/kWh) and corresponding total fleet vehicle-to-grid (V2G) revenue (€) under flat, peak, and solar pricing schemes for Chicago and San Francisco.
@@ 5.6 Sensitivity analysis of battery size and energy efficiency (San Francisco)
Vehicles 2 and 3 are more sensitive to battery size. For Vehicle 2, the median wait time decreases slightly from 225 s at 40 kWh to 217 s at 70 kWh, while the share of passengers waiting more than 600 s declines from 154 to 151. Increasing the battery to 100 kWh provides only marginal improvements, with the median wait time rising slightly to 219 s and the number of passengers waiting more than 600 s falling to 144. This suggests that a 70 kWh battery is sufficient to mitigate charging as a performance bottleneck for Vehicle 2, with further increases offering diminishing returns. For Vehicle 3, the least efficient option, the effect of battery size is far more pronounced. With 40 kWh, the system exhibits very poor performance, with a median wait time of 325s, a 95th percentile of 2625s, and more than 3600 passengers experiencing delays over 600s. Upgrading to 70 kWh brings dramatic improvements: the median drops to 219s, the 95th percentile to 432s, and delays over 600s fall by more than 95% to just 145. Increasing to 100 kWh does not significantly improve service further, with a median of 222s and 149 delays over 600s. This highlights that for inefficient vehicles, battery size is a critical driver of feasibility, with performance plateauing around 70 kWh.
@@
These findings are consistently reflected in other system-level metrics. Figures 9 demonstrate that increasing fleet size and charger availability effectively reduce wait times, particularly for the less efficient Vehicle 3 and some for Vehicle 2. Vehicle 1, however, continues to exhibit stable performance, characterized by lower 75th percentile and minimal sensitivity to infrastructure constraints. Furthermore, Vehicle 3 has the highest energy consumption and consequently the greatest charging demand, especially during peak hours, see Fig. 10(a). Conversely, Vehicle 1 maintains a smoother, more aligned charging-consumption profile throughout the day.
Fig.11(a) reports the estimated number of daily charge cycles (defined as charging from 20% to 80% state of charge) across vehicle types and battery sizes. As anticipated, larger batteries consistently require fewer charging cycles. Vehicle 1 consistently exhibits the lowest and most stable charging frequency, with a median of just under one cycle per day even with a 40 kWh battery. In contrast, Vehicles 2 and 3 demonstrate higher and more variable charging frequencies, particularly with smaller battery configurations. For instance, Vehicle 3 with a 40 kWh pack exhibits a median of just under 1.5 cycles per day and substantial variability in its charging behavior. Increasing the battery capacity to 70 kWh or 100 kWh nearly halves the median number of cycles and significantly narrows the distribution of charging events.
@@ 6 Conclusion
This paper introduced a combined stochastic and robust optimization framework for Electric Autonomous Mobility-on-Demand (EAMoD) operations under multiple uncertainties. By integrating Bayesian neural networks for demand and charger availability forecasting, clustering-based scenario tree reduction, and a tailored Nested Benders Decomposition (NBD) algorithm, we developed a scalable stochastic model predictive control (SMPC) approach capable of handling temporally correlated, multivariate uncertainties. High-fidelity simulations in San Francisco and Chicago demonstrate that SMPC reduces median passenger wait times by up to 36%, 95th percentile wait times by nearly 20%, rebalancing distance by up to 27%, and electricity costs by more than 35% relative to established baselines. These gains are consistent across two cities with very different demand intensities, spatial layouts, and fleet sizes, underscoring the framework’s robustness and transferability to diverse urban contexts.
Sensitivity analyses revealed that vehicle efficiency and battery size critically affect system performance. Energy-efficient vehicles maintain stable service levels even with small (40 kWh) batteries, requiring fewer charge cycles and exhibiting minimal sensitivity to infrastructure constraints. In contrast, less efficient vehicles show significant performance degradation with small batteries and benefit most from increases up to 70 kWh, beyond which improvements plateau, especially in high-demand settings such as Chicago. Moreover, efficiency not only drives service quality but also expands opportunities for energy-market participation: efficient vehicles consistently generate higher vehicle-to-grid (V2G) revenues, as their lower energy consumption leaves more capacity for grid interaction.
Finally, our results highlight that relying solely on median demand forecasts is insufficient, as deviations from the median can significantly affect operational outcomes. The use of probabilistic forecasts and scenario trees is therefore essential to capture predictive uncertainty and to design control strategies that remain robust under a wide range of future realizations. In this context, the tailored NBD algorithm proved crucial for solving large-scale scenario-based formulations efficiently, making SMPC computationally tractable in real-world urban settings. Taken together, the findings demonstrate the importance of jointly optimizing control strategies, vehicle selection, and battery configuration to ensure scalable, reliable, and sustainable EAMoD operations.